# Susceptibilities and Taylor coefficients of magnetic QCD from perturbation theory

### v1

@@ I Introduction
The phase diagram of strong interactions is ultimately to be constructed from in-medium quantum chromodynamics (QCD), the fundamental theory for hadronic matter. Besides temperature and different chemical potentials, the inclusion of an external magnetic field as one of the control parameters has proven to be phenomenologically relevant in different scenarios [1], from the astrophysics of compact stars [2, 3, 4] and binary neutron star mergers [5, 6] to non-central, high-energy heavy ion collisions [7, 8, 9, 10, 11, 12, 13], as well as in the early universe [14, 15, 16].
@@
First-principle calculations within magnetic QCD can be performed using Lattice QCD simulations [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28] (see Ref. [29] for a recent review), perturbation theory [30, 31, 32, 33, 34, 35], hard thermal loop perturbation theory [36, 37, 38, 39, 40], and in certain limits of QCD such as for a large number of colors $N_{c}$ [41] and within chiral perturbation theory [42, 43, 44, 45]. Of course, one can also adopt holographic models [46, 47, 48, 49, 50] and a variety of effective models [51, 52, 53, 54, 55, 1].
@@
Recently, the QCD equation of state in the presence of an external magnetic field $B$ and nonzero baryon density $\mu_{B}$, besides the temperature $T$, was investigated via lattice simulations with $2+1$ dynamical staggered quarks at their physical masses [56]. The simulations were performed using the method of the imaginary baryon chemical potential to circumvent the Sign Problem [57]. Their results indicate considerable enhancement of the coefficients for the expansion of the pressure in powers of $\mu_{B}/T$ by the presence of the magnetic field, considering values up to $eB=1.2$ GeV2. The presence of the magnetic field also shifts their characteristic temperature dependence to lower values, which is consistent with the reduction of the critical temperature for the chiral transition and a strengthening of the transition as the magnetic field is increased [29].
@@
In this paper we compute the coefficients $c_{2}(T,B)$ and $c_{4}(T,B)$ for the expansion of the pressure in powers of $\mu_{B}/T$ in the presence of a large magnetic field from first principles within perturbative QCD (pQCD) at finite temperature and baryon density up to two-loops for $N_{f}=3$ flavors with physical quark masses. Since we use perturbative QCD within the lowest-Landau level (LLL) approximation, the region of validity for our framework is given by $m_{s}\ll T\ll\sqrt{eB}$, where $m_{s}$ is the strange quark mass, $e$ is the fundamental electric charge, and $B$ is the magnetic field strength. We include the effects of the renormalization scale in the running coupling, $\alpha_{s}(T,\sqrt{eB})$, and running quark masses, since these effects have proven to be relevant for the resulting thermodynamics [58, 59, 60, 61, 62].
@@ II Theoretical framework
Let us start with the pressure of thermal QCD in the presence of high magnetic fields, which has been computed to two-loop order [30, 33, 34]. The one-loop (free), contribution from the quark sector is given by the following renormalized expression (subtracting the pure vacuum term) [51, 52, 53, 54]:
$\begin{split}\frac{P_{\rm free}^{q}}{N_{c}}=&\sum_{f}\frac{(q_{f}B)^{2}}{2\pi^{2}}\left[\zeta^{\prime}(-1,x_{f})-\zeta^{\prime}(-1,0)+\frac{1}{2}\left(x_{f}-x_{f}^{2}\right)\ln x_{f}+\frac{x_{f}^{2}}{2}\right]\\ &+T\sum_{n,f}\frac{q_{f}B}{\pi}\left(1-\delta_{n,0}/2\right)\int\frac{dp_{z}}{2\pi}\bigg{\{}\ln\left(1+e^{-\beta\left[E(n,p_{z})-\mu_{f}\right]}\right)+\ln\left(1+e^{-\beta\left[E(n,p_{z})+\mu_{f}\right]}\right)\bigg{\}}\,,\end{split}$ (1)
where $E^{2}(n,p_{z})=p_{z}^{2}+m_{f}^{2}+2q_{f}Bn$, $x_{f}\equiv m_{f}^{2}/2q_{f}B$, $T=1/\beta$ is the temperature, $\mu_{f}$ is the quark chemical potential, $N_{c}$ is the number of colors, $f$ labels quark flavors, $q_{f}$ is the quark electric charge, and $n=0,1,2,\cdots$ stands for the Landau levels. In this expression, Matsubara sums have already been performed in the medium contribution [^1] One should notice that there is an inherent arbitrariness in the renormalization procedure (see Refs.[63, 64, 65, 66, 67, 68, 69, 70, 71] for a discussion). In Eq. (1), all mass-independent terms were neglected and the pure magnetic term goes to zero in the limit $m\to 0$. There are renormalization procedures where other terms survive and the pure magnetic expression diverges as $m\to 0$. This discrepancy in the renormalized expression leads to differences in some physical quantities, e.g. the magnetization [66]..
@@
The one-loop contribution from the gluons has the usual Stefan-Boltzmann form [72]
@@
Here $\bm{k}_{L}=(i\omega_{\ell}^{B},k_{z})$, $\bm{p}_{L}=(i\omega_{n_{1}}^{F}+\mu_{f},p_{z})$, $\bm{q}_{L}=(i\omega_{n_{2}}^{F}+\mu_{f},q_{z})$.
Evaluating the momentum integrations yields [34]
$\begin{split}\mathcal{G}(m_{k}^{2},m_{f}^{2})=-\beta Vg^{2}\left(\frac{N_{c}^{2}-1}{2}\right)T^{2}m_{f}^{2}\sum_{\ell,n_{2}}\frac{\mathcal{E}_{\ell}-\mathcal{E}_{n_{2}}}{\mathcal{E}_{\ell}\mathcal{E}_{n_{1}}\mathcal{E}_{n_{2}}\left|\mathcal{E}_{\ell}-\mathcal{E}_{n_{2}}\right|\left(\left|\mathcal{E}_{\ell}-\mathcal{E}_{n_{2}}\right|+\mathcal{E}_{n_{1}}\right)}\,,\end{split}$ (6)
@@
The two-loop contribution from the gluons is given by the well-known formula [72]:
@@
Those can be directly compared to recent lattice results [56], as we show in the sequel.
@@
The running of both $\alpha_{s}$ and $m_{s}$ are known to four-loop order in the $\overline{\rm MS}$ scheme [73]. For our $O(\alpha_{s})$ calculation of the QCD pressure, we adopt for the coupling [60]
@@
[…]
with $\hat{m}_{s}$ being the renormalization group invariant strange quark mass, i.e. $\bar{\Lambda}$ independent. Since Eq. (14) for $\alpha_{s}$ tells us that different values of $N_{f}$ give different values of $\Lambda_{\overline{\rm MS}}$, by choosing $\alpha_{s}(\bar{\Lambda}=1.5~{}{\rm GeV},~{}N_{f}=3)=0.336^{+0.012}_{-0.008}$ [74], we obtain $\Lambda^{2+1}_{\overline{\rm MS}}=343^{+18}_{-12}~{}$eV. Finally, fixing the strange quark mass at $m_{s}(2~{}{\rm GeV},N_{f}=3)=92.4(1.5)~{}$MeV [75] gives $\hat{m}^{2+1}_{s}~{}{\approx}~{}248.7~{}$MeV when using $\alpha^{2+1}_{s}$ in Eq. (15).
Following the detailed analysis of different implementations of the scale dependence performed in Ref. [33], we show results only for the most physical case: $\bar{\Lambda}=\sqrt{(2\pi T)^{2}+eB}$, which corresponds to a natural extension of what is done in finite-temperature field theory in the absence of a magnetic field [72]. Hence, we show results from pQCD including bands that encode the renormalization-scale dependence in the standard range between half the central scale $\bar{\Lambda}$ and twice its value.
@@ III Results
Figure 1: Coefficient $c_{2}$ as a function of the temperature for $eB=1.2$ GeV2. Lattice data (lower, red band) from Ref. [56] and bands on the right from pQCD up to one and two loops. Bands in the perturbative part represent the renormalization-scale dependence in the standard range between half the central scale $\bar{\Lambda}$ and twice its value.
In Figure 1 the coefficient $c_{2}$ as a function of the temperature is displayed for $eB=1.2$ GeV2, comparing our pQCD results with the lattice data from Ref. [56]. In this reference, lattice simulations are performed for chemical potentials such that $\mu_{u}=\mu_{d}$ and $\mu_{s}=0$. In our perturbative calculations that corresponds to having no contribution from the $s$ quark to the coefficients $c_{n}$. That is precisely the reason why the one-loop contribution, $c_{2}^{\rm{1L}}$, has no scale dependence, whereas the full two-loop coefficient, $c_{2}^{2L}$, displays a tiny theoretical uncertainty band: for $u$ and $d$ quarks, only the two-loop contribution has scale dependence and it is very small for small masses, since $c_{n}\sim m^{2}$, as can be seen in Eq. (6). Figure 1 clearly shows that the perturbative results seem to be compatible with those from lattice QCD for higher temperatures.
Figure 2: Coefficient $c_{4}$ as a function of the temperature for $eB=1.2$ GeV2. Lattice data (left band) from Ref. [56] and bands on the right from pQCD up to one and two loops. Bands in the perturbative part represent the renormalization-scale dependence in the standard range between half the central scale $\bar{\Lambda}$ and twice its value.
Figure 2 displays our results for the coefficient $c_{4}$ as a function of the temperature for $eB=1.2$ GeV2, as well as Lattice data from Ref. [56]. One can see that the lattice band merges quite smoothly onto the (narrow) perturbative band, so that there is a nice overlap region. The perturbative band at two-loop order is narrow, and the dislocation between one- and two-loop contributions is considerably larger [^2] This is highly dependent on the mass value. When there are contributions from the strange quark, the exchange contribution is larger, as can be seen in Fig. 7. than in the case of $c_{2}$. However, it still requires magnification of the plot to be visible and remains much smaller than the typical size of the lattice QCD band.
@@
Our predictions for the behavior of $c_{2}$ as function of the magnetic field for $T=500$ MeV are shown in Figure 3, going to much higher values of the magnetic field strength, up to $10$ GeV2. As the magnetic field increases, $c_{2}$ increases in a linear fashion with a very narrow perturbative band. The linear trend is in line with lattice data shown in Ref. [56] for three values of the magnetic field, $eB=1.2$ GeV2 being the highest. Our perturbative prediction then states that the linear behavior should extend also to much higher magnetic field values. The coefficient $c_{4}$ for large magnetic fields is, on the other hand, very small, essentially zero.
@@
Figure 4: The excess of pressure, $\Delta P$, as function of $\hat{\mu}_{B}$ for two values of temperature, $T=152,160$ MeV. Here, $eB=1.2$ $\rm{GeV}^{2}$. Lattice data from Ref. [56]. Bands in the perturbative results represent the renormalization-scale dependence in the standard range between half the central scale $\bar{\Lambda}$ and twice its value.
Figure 5: Baryon density as a function of $\hat{\mu}_{B}$ for two values of temperature, $T=152,160$ MeV. Here, $eB=1.2$ $\rm{GeV}^{2}$. Lattice data from Ref. [56]. Bands in the perturbative results represent the renormalization-scale dependence in the standard range between half the central scale $\bar{\Lambda}$ and twice its value.
Figure 6: Baryon number susceptibility as a function of $\hat{\mu}_{B}$ for two values of temperature, $T=152,160$ MeV. Here, $eB=1.2$ $\rm{GeV}^{2}$. Lattice data from Ref. [56]. Bands in the perturbative results represent the renormalization-scale dependence in the standard range between half the central scale $\bar{\Lambda}$ and twice its value.
@@
Finally, we present results for the susceptibilities in the $\left\{\mu_{B},\mu_{Q},\mu_{S}\right\}$ basis, which can be compared to the findings of Ref. [76] on the lattice for a smaller value of the magnetic field. There they are given by the following relations:
@@
Our perturbative two-loop calculation is diagonal in flavor, since non-diagonal contributions appear only at the next perturbative order due to the ring resummation [77]. Thus, we have $\chi_{110}=\chi_{101}=\chi_{011}=0$. In Figure 7 we show the susceptibilities as functions of the temperature for a magnetic field $eB=0.8$ $\rm{GeV}^{2}$. We compare our perturbative results with lattice data from reference [76]. Here, we can notice that the strange contribution has a larger scale dependence due to the contribution of $m_{s}$, as discussed before. Even though the magnetic field is probably too low for our LLL hierarchy of scales to hold, it is encouraging that our results are in the same ballpark of Lattice QCD for $T\gtrsim 170$ MeV.
Figure 7: $\chi_{BB}$, $\chi_{QQ}$ and $\chi_{SS}$ as functions of the temperature for $eB=0.8$ $\rm{GeV}^{2}$. Lattice data from Ref. [76]. Bands in the perturbative results represent the renormalization-scale dependence in the standard range between half the central scale $\bar{\Lambda}$ and twice its value.

### v2

@@ I Introduction
The phase diagram of strong interactions is ultimately to be constructed from in-medium quantum chromodynamics (QCD), the fundamental theory for hadronic matter. Besides temperature and different chemical potentials, the inclusion of an external magnetic field as one of the control parameters has proven to be phenomenologically relevant in different scenarios Adhikari et al. [2024], from the astrophysics of compact stars Duncan and Thompson [1992], Thompson and Duncan [1993], Kouveliotou et al. [1998] and binary neutron star mergers Jiang et al. [2025], Most et al. [2025] to non-central, high-energy heavy ion collisions Kharzeev et al. [2008], Skokov et al. [2009], Voronyuk et al. [2011], Bzdak and Skokov [2012], Deng and Huang [2012], Inghirami et al. [2016], Roy et al. [2017], as well as in the early universe Vachaspati [1991], Enqvist and Olesen [1993], Grasso and Rubinstein [2001].
@@
First-principle calculations within magnetic QCD can be performed using Lattice QCD simulations D’Elia et al. [2010], Bali et al. [2012a], Ilgenfritz et al. [2012], Bali et al. [2012b], Bornyakov et al. [2014], Bali et al. [2013], Bruckmann et al. [2013], Bali et al. [2014], Endrődi [2015], D’Elia et al. [2018, 2021, 2022] (see Ref. Endrődi [2024] for a recent review), perturbation theory Blaizot et al. [2013], Ayala et al. [2015, 2016], Fraga et al. [2023, 2024a, 2024b], hard thermal loop perturbation theory Rath and Patra [2017], Haque [2017], Karmakar et al. [2019], Bandyopadhyay et al. [2019], Karmakar et al. [2020], and in certain limits of QCD such as for a large number of colors $N_{c}$ Fraga et al. [2013] and within chiral perturbation theory Colucci et al. [2014], Hofmann [2020a, b], Adhikari and Andersen [2021]. Of course, one can also adopt holographic models Preis et al. [2011, 2012, 2013], Finazzo et al. [2016], Critelli et al. [2016] and a variety of effective models Fraga [2013], Kharzeev et al. [2013], Andersen et al. [2016], Miransky and Shovkovy [2015], Andersen [2021], Adhikari et al. [2024].
@@
Recently, the QCD equation of state in the presence of an external magnetic field $B$ and nonzero baryon density $\mu_{B}$, besides the temperature $T$, was investigated via lattice simulations with $2+1$ dynamical staggered quarks at their physical masses Astrakhantsev et al. [2024]. The simulations were performed using the method of the imaginary baryon chemical potential to circumvent the Sign Problem Aarts [2016]. Their results indicate considerable enhancement of the coefficients for the expansion of the pressure in powers of $\mu_{B}/T$ by the presence of the magnetic field, considering values up to $eB=1.2$ GeV2. The presence of the magnetic field also shifts their characteristic temperature dependence to lower values, which is consistent with the reduction of the critical temperature for the chiral transition and a strengthening of the transition as the magnetic field is increased Endrődi [2024].
@@
In this paper we compute the coefficients $c_{2}(T,B)$ and $c_{4}(T,B)$ for the expansion of the pressure in powers of $\mu_{B}/T$ in the presence of a large magnetic field from first principles within perturbative QCD (pQCD) at finite temperature and baryon density up to two-loops for $N_{f}=3$ flavors with physical quark masses. Since we use perturbative QCD within the lowest-Landau level (LLL) approximation, the region of validity for our framework is given by $m_{s}\ll T\ll\sqrt{eB}$, where $m_{s}$ is the strange quark mass, $e$ is the fundamental electric charge, and $B$ is the magnetic field strength. We include the effects of the renormalization scale in the running coupling, $\alpha_{s}(T,\sqrt{eB})$, and running quark masses, since these effects have proven to be relevant for the resulting thermodynamics Freedman and McLerran [1978], Farhi and Jaffe [1984], Fraga and Romatschke [2005], Kurkela et al. [2010], Gorda and Säppi [2022].
@@ II Theoretical framework
Let us start with the pressure of thermal QCD in the presence of high magnetic fields, which has been computed to two-loop order Blaizot et al. [2013], Fraga et al. [2023, 2024a]. The one-loop (free), contribution from the quark sector is given by the following renormalized expression (subtracting the pure vacuum term) Fraga [2013], Kharzeev et al. [2013], Andersen et al. [2016], Miransky and Shovkovy [2015]:
[…]
where $E^{2}(n,p_{z})=p_{z}^{2}+m_{f}^{2}+2q_{f}Bn$, $x_{f}\equiv m_{f}^{2}/2q_{f}B$, $T=1/\beta$ is the temperature, $\mu_{f}$ is the quark chemical potential, $N_{c}$ is the number of colors, $f$ labels quark flavors, $q_{f}$ is the quark electric charge, and $n=0,1,2,\cdots$ stands for the Landau levels. In this expression, Matsubara sums have already been performed in the medium contribution [^1] One should notice that there is an inherent arbitrariness in the renormalization procedure in the presence of a magnetic background (see Refs. Fraga and Mizher [2008], Mizher et al. [2010], Fraga and Palhares [2012], Endrődi [2013], Haber et al. [2014], Avancini et al. [2019, 2021], Tavares et al. [2021], Farias et al. [2022] for a discussion). In Eq. (1), all mass-independent terms were neglected and the pure magnetic term goes to zero in the limit $m\to 0$. There are renormalization procedures where other terms survive and the pure magnetic expression diverges as $m\to 0$. This discrepancy in the renormalized expression leads to differences in some physical quantities, e.g. the magnetization Endrődi [2013]. It is important to note however, that the susceptibilities and the expansion coefficients calculated below are independent of the renormalization procedure, since the pure magnetic terms do not contribute..
@@
The one-loop contribution from the gluons has the usual Stefan-Boltzmann form Kapusta and Gale [2011]
@@
Here $\bm{k}_{L}=(i\omega_{\ell}^{B},k_{z})$, $\bm{p}_{L}=(i\omega_{n_{1}}^{F}+\mu_{f},p_{z})$, $\bm{q}_{L}=(i\omega_{n_{2}}^{F}+\mu_{f},q_{z})$, where $\omega_{\ell}^{B}=2\pi\ell T$ for bosons and $\omega_{n_{i}}^{F}=(2n_{i}+1)\pi T$ for fermions are Matsubara frequencies, with $\ell$ and $n_{i}$ integers.
Evaluating the momentum integrations yields  [^2] In this expression, the UV limit is controlled by the Gaussian weight $e^{-\frac{m_{k}^{2}}{2q_{f}B}}$, while the IR sector is controlled by the $m^{2}_{f}$ overall factor. As a result, the two-loop pressure in the presence of an extremely strong magnetic background vanishes in the chiral limit Blaizot et al. [2013], Fraga et al. [2024a].
[…]
@@
As remarked long ago Blaizot et al. [2013], the chiral limit for perturbative QCD at very high magnetic fields is extremely simple: the exchange contribution vanishes identically for $m_{f}=0$ (see also Ref. Fraga et al. [2023]).
The two-loop contribution from the gluons is given by the well-known formula Kapusta and Gale [2011]:
@@
Those can be directly compared to recent lattice results Astrakhantsev et al. [2024], as we show in the sequel.
@@
The running of both $\alpha_{s}$ and $m_{s}$ are known to four-loop order in the $\overline{\rm MS}$ scheme Vermaseren et al. [1997]. For our $O(\alpha_{s})$ calculation of the QCD pressure, we adopt for the coupling Fraga and Romatschke [2005]
@@
$\begin{aligned} m_{s}(\bar{\Lambda})=\hat{m}_{s}\left(\frac{\alpha_{s}}{\pi}\right)^{4/9}\left[1+0.895062\left(\frac{\alpha_{s}}{\pi}\right)\right]\;,\end{aligned}$ (15)
with $\hat{m}_{s}$ being the renormalization group invariant strange quark mass, i.e. $\bar{\Lambda}$ independent. Since Eq. (14) for $\alpha_{s}$ tells us that different values of $N_{f}$ give different values of $\Lambda_{\overline{\rm MS}}$, by choosing $\alpha_{s}(\bar{\Lambda}=1.5~{\rm GeV},~N_{f}=3)=0.336^{+0.012}_{-0.008}$ Bazavov et al. [2014], we obtain $\Lambda^{2+1}_{\overline{\rm MS}}=343^{+18}_{-12}$~MeV. Finally, fixing the strange quark mass at $m_{s}(2~{\rm GeV},N_{f}=3)=92.4(1.5)$~MeV Chakraborty et al. [2015] gives $\hat{m}^{2+1}_{s}~{\approx}~248.7$~MeV when using $\alpha^{2+1}_{s}$ in Eq. (15).
Following the detailed analysis of different implementations of the scale dependence performed in Ref. Fraga et al. [2023], we show results only for the most physical case: $\bar{\Lambda}=\sqrt{(2\pi T)^{2}+eB}$, which corresponds to a natural extension of what is done in finite-temperature field theory in the absence of a magnetic field Kapusta and Gale [2011]. Hence, we show results from pQCD including bands that encode the renormalization-scale dependence in the standard range between half the central scale $\bar{\Lambda}$ and twice its value.
@@ III Results
Figure 1: Coefficient $c_{2}$ as a function of the temperature for $eB=1.2$ GeV2. Lattice data (lower, red band) from Ref. Astrakhantsev et al. [2024] and bands on the right from pQCD up to one and two loops. Bands in the perturbative part represent the renormalization-scale dependence in the standard range between half the central scale $\bar{\Lambda}$ and twice its value.
In Figure 1 the coefficient $c_{2}$ as a function of the temperature is displayed for $eB=1.2$ GeV2, comparing our pQCD results with the lattice data from Ref. Astrakhantsev et al. [2024]. In this reference, lattice simulations are performed for chemical potentials such that $\mu_{u}=\mu_{d}$ and $\mu_{s}=0$. In our perturbative calculations that corresponds to having no contribution from the $s$ quark to the coefficients $c_{n}$. That is precisely the reason why the one-loop contribution, $c_{2}^{\rm{1L}}$, has no scale dependence, whereas the full two-loop coefficient, $c_{2}^{2L}$, displays a tiny theoretical uncertainty band: for $u$ and $d$ quarks, only the two-loop contribution has scale dependence and it is very small for small masses, since $c_{n}\sim m^{2}$, as can be seen in Eq. (6). Figure 1 clearly shows that the perturbative results seem to be compatible with those from lattice QCD for higher temperatures.
Figure 2: Coefficient $c_{4}$ as a function of the temperature for $eB=1.2$ GeV2. Lattice data (left band) from Ref. Astrakhantsev et al. [2024] and bands on the right from pQCD up to one and two loops. Bands in the perturbative part represent the renormalization-scale dependence in the standard range between half the central scale $\bar{\Lambda}$ and twice its value.
Figure 2 displays our results for the coefficient $c_{4}$ as a function of the temperature for $eB=1.2$ GeV2, as well as Lattice data from Ref. Astrakhantsev et al. [2024]. One can see that the lattice band merges quite smoothly onto the (narrow) perturbative band, so that there is a nice overlap region. The perturbative band at two-loop order is narrow, and the dislocation between one- and two-loop contributions is considerably larger [^3] This is highly dependent on the mass value. When there are contributions from the strange quark, the exchange contribution is larger, as can be seen in Fig. 7. than in the case of $c_{2}$. However, it still requires magnification of the plot to be visible and remains much smaller than the typical size of the lattice QCD band.
@@
Our predictions for the behavior of $c_{2}$ as function of the magnetic field for $T=500$ MeV are shown in Figure 3, going to much higher values of the magnetic field strength, up to $10$ GeV2. As the magnetic field increases, $c_{2}$ increases in a linear fashion with a very narrow perturbative band. The linear trend is in line with lattice data shown in Ref. Astrakhantsev et al. [2024] for three values of the magnetic field, $eB=1.2$ GeV2 being the highest. Our perturbative prediction then states that the linear behavior should extend also to much higher magnetic field values. The coefficient $c_{4}$ for large magnetic fields is, on the other hand, very small, essentially zero.
@@
Figure 4: The excess of pressure, $\Delta P$, as function of $\hat{\mu}_{B}$ for two values of temperature, $T=152,160$ MeV. Here, $eB=1.2$ $\rm{GeV}^{2}$. Lattice data from Ref. Astrakhantsev et al. [2024]. Bands in the perturbative results represent the renormalization-scale dependence in the standard range between half the central scale $\bar{\Lambda}$ and twice its value.
Figure 5: Baryon density as a function of $\hat{\mu}_{B}$ for two values of temperature, $T=152,160$ MeV. Here, $eB=1.2$ $\rm{GeV}^{2}$. Lattice data from Ref. Astrakhantsev et al. [2024]. Bands in the perturbative results represent the renormalization-scale dependence in the standard range between half the central scale $\bar{\Lambda}$ and twice its value.
Figure 6: Baryon number susceptibility as a function of $\hat{\mu}_{B}$ for two values of temperature, $T=152,160$ MeV. Here, $eB=1.2$ $\rm{GeV}^{2}$. Lattice data from Ref. Astrakhantsev et al. [2024]. Bands in the perturbative results represent the renormalization-scale dependence in the standard range between half the central scale $\bar{\Lambda}$ and twice its value.
@@
Finally, we present results for the susceptibilities in the $\left\{\mu_{B},\mu_{Q},\mu_{S}\right\}$ basis, which can be compared to the findings of Ref. Borsanyi et al. [2024] on the lattice for a smaller value of the magnetic field. There they are given by the following relations:
@@
Our perturbative two-loop calculation is diagonal in flavor, since non-diagonal contributions appear only at the next perturbative order due to the ring resummation Freedman and McLerran [1977]. Thus, we have $\chi_{110}=\chi_{101}=\chi_{011}=0$. In Figure 7 we show the susceptibilities as functions of the temperature for a magnetic field $eB=0.8$ $\rm{GeV}^{2}$. We compare our perturbative results with lattice data from reference Borsanyi et al. [2024]. Here, we can notice that the strange contribution has a larger scale dependence due to the contribution of $m_{s}$, as discussed before. Even though the magnetic field is probably too low for our LLL hierarchy of scales to hold, it is encouraging that our results are in the same ballpark of Lattice QCD for $T\gtrsim 170$ MeV.
Figure 7: $\chi_{BB}$, $\chi_{QQ}$ and $\chi_{SS}$ as functions of the temperature for $eB=0.8$ $\rm{GeV}^{2}$. Lattice data from Ref. Borsanyi et al. [2024]. Bands in the perturbative results represent the renormalization-scale dependence in the standard range between half the central scale $\bar{\Lambda}$ and twice its value.
It is clear that lattice simulations are not within the ideal hierarchy of scales to provide a fair comparison to perturbative QCD yet. In this regard, it would be necessary to have more values of the magnetic field from the lattice to compare in the future. Nevertheless, our results are in a reasonably good agreement, which is encouraging. They, of course, miss the non-monotonic behavior of the coefficients $c_{2}$ and $c_{4}$, as expected from a perturbative calculation. On the other hand, the reduction of the contribution from higher orders in perturbation theory and the narrow uncertainty bands from the renormalization scale freedom make the description reasonable.